9:24: system fully started, LGSW, ADSEC, FLAO, IRTC are ready. On the FLAO we load the board setup 2014-05-16 - ARGOS cal onaxis.

Board stage mm
X -70
Y -49
Z 37
We begin creating different subapertures sets for different slope computing algorithms. The parent subaperture set is 20141104_192000, the subapertures positions and numbers are not changed, just the pixel weights are variated.

SL. ALGO SUBAPS Notes
CoG 20141104_192000 Standard algo, px of each subap are normalized on the subap flux
WWPA 20141106_162533 px value multiplied by the sqrt of the px flux
CoG Norm 20141106_163503 all px are normalized on the overall pupil average flux
ToCG48 20141106_180427 subaps thresholded to 48 ADU

We reposition hexapod to send light into LGSW. We close the jitter loop we have these values:

Pupil PI X PI Y
Blue -3.40 -0.15
Yellow -3.80 4.60
Red -2.30 -3.45
HEXAPOD: X = -3.02, Y = -6.27, Z = 4.38, Rx = -244, Ry = 27

Now we close the loop with few modes to better center the hexapod on the 3 FoVs:

Pupil PI X PI Y
Blue 1.00 0.00
Yellow -1.80 0.70
Red 0.00 0.45
HEXAPOD: X = -3.02, Y = -6.27, Z = 4.38, Rx = -244, Ry = 30

We close the loop on the 150modes without disturbance, in the meanwhile we build a reconstructor with 50modes.

ARGOS TN FLAO TN Notes Res.
20141106_172652   CL, g= [0.5, 0.4] 80
20141106_173035   " 103
20141106_172652-20141106_173035_modalplot.jpg

REC: 50modes_ryb_300Hz_TTremoved

We load the new reconstructor, we set the diodes to the minimum current (0.03A = 500-1000 ADU/subap). We close the loop with gain [0.5, 0.3] and also the pupil loop is closed. Everything seems OK. The blue pupil goes away but it recovers by itself...

ARGOS TN FLAO TN Notes Res.
20141106_175144   New 50modes REC, HIGH FLUX, CL, g= [0.5, 0.4] 70
20141106_175655   New 50modes REC, LOW FLUX, CL, g= [0.5, 0.4] 162
20141106_175745   " 162
20141106_175144-20141106_175745_modalplot.jpg

PBL controller crashes because it lost the connection with BCU. Anyhow the loop is still running...

Acquisition of IMs for different slope computation algorithms

We power up the diodes to have a flux of about 8000 ADU/subap on each pupil.

11:20: we have to restart the controller because it lost the connection with BCU. Then we have to recenter hexapod and PIs:

Pupil PI X PI Y
Blue -1.67 -0.67
Yellow -1.00 0.73
Red -0.83 0.45
HEXAPOD: X = -3.02, Y = -6.27, Z = 4.38, Rx = -247, Ry = 32

We start calibrating the various slope algorithms, the pupil loop is closed (PCL) and the jitter loop is open (JOL). The acquisition sequence is the same for all the algos:
  1. acquire IM in OL
  2. generate a 10 modes REC, close the loop with it and acquire another IM
  3. generate a 50 modes REC, close the loop with it and acquire another IM
  4. generate a 150 modes REC

Algo IM REC Notes
TCoG48 20141106_183040 10modes_ryb_1kHz_TTrm_Thres48  
20141106_183808 50modes_ryb_1kHz_TTrm_Thres48  
20141106_184216 150modes_ryb_1kHz_TTrm_Thres48  
CoG Norm 20141106_184613 wrong  
20141106_184613 10modes_ryb_1kHz_TTrm_CoGNorm  
20141106_185448 50modes_ryb_1kHz_TTrm_CoGNorm  
20141106_185902 150modes_ryb_1kHz_TTrm_CoGNorm  
WWPA 20141106_190503 10modes_ryb_1kHz_TTrm_WWPA  
20141106_190927 50modes_ryb_1kHz_TTrm_WWPA  
20141106_191350 150modes_ryb_1kHz_TTrm_WWPA  
CoG 20141106_192014 10modes_ryb_1kHz_TTrm_CoG  
20141106_192431 50modes_ryb_1kHz_TTrm_CoG  
20141106_192845 150modes_ryb_1kHz_TTrm_CoG  

Acquisition of slope offset

Now we optimize the PSF on CCD47 by closing the FLAO loop with 400 modes, we use FLAO REC: 20130919_014135, we put the white light source at 80 steps. At this point we acquire the slope offset for all the different slope computation algorithms, in the meanwhile we acquire snapshots on both ARGOS and FLAO:

Algo Slope Offset ARGOS TN FLAO TN Res.
CoG 20141106_195056 20141106_195214 20141106_194936 46
20141106_195236   46
WWPA 20141106_195643 20141106_195658 20141106_195535 37
20141106_195724   40
CoGNorm 20141106_195833 20141106_195839 20141106_195721 42
20141106_195900   42
TCoG48 20141106_200019 20141106_200032 20141106_195909 33
20141106_200058   33
SlopeAlgo_comparison_FLAO_CL.jpg

High flux test

We close now the loop on ARGOS, PCL, JOL, we load the 0.8" disturbance.

Algo ARGOS TN FLAO TN Res.
OL 20141106_200903 20141106_200741 466
20141106_200923   465
20141106_200945   465
CoG 20141106_200720 20141106_200554 59
20141106_200744 20141106_200639 59
WWPA 20141106_201451 20141106_201300 37
20141106_201516 20141106_201419 37
20141106_201540   37
CoGNorm 20141106_202117 20141106_201959 54
20141106_202144 20141106_202128 50
20141106_202202   51
TCoG48 20141106_203501 20141106_203337 38
20141106_203521 20141106_203456 35
20141106_203539   38
SlopeAlgo_comparison_ARGOS_CL_8k.jpg

Low flux test

Algo ARGOS TN FLAO TN Res. Notes
CoG 20141106_205751 20141106_205652 160 g=[0.4,0.3,0.1], POL
20141106_205813   160 "
WWPA 20141106_210340   94 g=[0.5,0.4,0.2], POL
20141106_210405 20141106_210222 1117 Loop fails!
20141106_210603 20141106_210442 101 g=[0.4,0.3,0.1]
20141106_210625   101 same gain, PCL
CoG 20141106_211836 20141106_211714 183 g=[0.5,0.3,0.0], POL
20141106_211858   168 "
TCoG48 20141106_212356   87 g=[0.5,0.3,0.2]
20141106_212433 20141106_212233 90 g=[0.5,0.4,0.3]
20141106_212456   89 "
20141106_212531   89 "
20141106_212715 20141106_212550 97 g=[0.5,0.4,0.3], PCL
20141106_212739   95 "
20141106_212818 20141106_212649 95 "
WWPA 20141106_213343 20141106_213220 98 g=[0.5,0.4,0.3]
20141106_213405 20141106_213304 98 "
SlopeAlgo_comparison_ARGOS_CL_1k.jpg

14:30 : we have repeated problems with BCU: it becomes unreachable and we have to reset it (it happens 3 times in a row)
Topic revision: r3 - 22 Nov 2014, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback